Show a relationship between potential, kinetic and rest energy proving which is normally higher than the rest
1
Expert's answer
2019-09-24T12:07:32-0400

When the mass is moving then the moving mass posses more energy then the mass at rest due to Lorentz contraction factor.

Moving mass energy is the sum of relativistic kinetic energy and rest mass energy. Thus,

Relativistic kinetic energy is the difference of total energy due to moving mass and the rest mass energy.

Relativistic kinetic energy is KErel = (γ − 1)mc2

"\gamma=\frac{1}{\sqrt{ (1-v^2/c^2) }}"